OpenStudy (anonymous):

In a 30-60-90 triangle if the short leg is 3 inches, what is the measure of the long leg and the hypotenuse?

OpenStudy (anonymous):

the long leg is always \[\sqrt{3}\] and the hypotenus is always double the short leg

OpenStudy (anonymous):

Okay, so 60√3?

OpenStudy (anonymous):

\[3*\sqrt{3}\]

OpenStudy (anonymous):

3 can't go into \[\sqrt{3}\] so it will be \[3\sqrt{3}\]

OpenStudy (anonymous):

Okay so, then it turns into √6?

OpenStudy (anonymous):

Wait, 9 XD

OpenStudy (anonymous):

It stays like that because whole numbers and numbers under the radical don't combine

OpenStudy (anonymous):

Oh okay, so then is that the long leg or the hypotenuse?

OpenStudy (anonymous):

long leg

OpenStudy (anonymous):

Wait, so it's the long leg, and the hypotenuse is double that right?

OpenStudy (anonymous):

the hypotenus is double the short leg

OpenStudy (anonymous):

so 3*2

OpenStudy (anonymous):

Okay! Thanks :) Can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

In a 45-45-90 triangle, if one leg is 10ft, what is the measure of the other leg and the hypotenuse?

OpenStudy (anonymous):

in this triangle the legs are x, and the hypotenuse is \[x \sqrt{2}\]